# SimGRAG: Leveraging Similar Subgraphs for Knowledge Graphs Driven Retrieval-Augmented Generation

### v1

@@ Abstract
Recent advancements in large language models (LLMs) have shown impressive versatility across various tasks. To eliminate its hallucinations, retrieval-augmented generation (RAG) has emerged as a powerful approach, leveraging external knowledge sources like knowledge graphs (KGs). In this paper, we study the task of KG-driven RAG and propose a novel Similar Graph Enhanced Retrieval-Augmented Generation (SimGRAG) method. It effectively addresses the challenge of aligning query texts and KG structures through a two-stage process: (1) query-to-pattern, which uses an LLM to transform queries into a desired graph pattern, and (2) pattern-to-subgraph, which quantifies the alignment between the pattern and candidate subgraphs using a graph semantic distance (GSD) metric. We also develop an optimized retrieval algorithm that efficiently identifies the top-$k$ subgraphs within 1-second latency on a 10-million-scale KG. Extensive experiments show that SimGRAG outperforms state-of-the-art KG-driven RAG methods in both question answering and fact verification, offering superior plug-and-play usability and scalability. Our code is available at https://github.com/YZ-Cai/SimGRAG.
@@ 1 Introduction
Pre-trained large language models (LLMs) have become popular for diverse applications due to their generality and flexibility Zhao et al. (2023); Minaee et al. (2024); Wang et al. (2024a). To avoid the hallucinations or outdated internal knowledge of LLMs Zhang et al. (2023); Baek et al. (2023), Retrieval-Augmented Generation (RAG) Zhao et al. (2024); Gao et al. (2023) integrates LLMs with external databases or documents to ensure the generated outputs are grounded in the most relevant and up-to-date information. Recently, knowledge graphs (KGs) have emerged as a valuable data source for RAG Peng et al. (2024), which encode knowledge through interconnected entities and relationships Ji et al. (2022). State-of-the-art KG-driven RAG methods, such as KAPING Baek et al. (2023), KG-GPT Kim et al. (2023a), KELP Liu et al. (2024), and G-retriever He et al. (2024), typically follow a paradigm of retrieving subgraphs from the KG and feeding them into LLMs to generate the final response. As shown in Figure 1, beyond achieving satisfactory answer quality, a practically ideal approach should address the following features.
@@
Plug-and-Play. To fully leverage the inherent generalization power of LLMs, an ideal approach should enable seamless integration with diverse KGs and LLMs without additional training or fine-tuning. Otherwise, training a smaller, task-specific model would be a more cost-effective alternative, as confirmed by our experiments in Section 6.3.
Avoidance of Entity Leaks. In large-scale KGs, numerous entities may somehow related to a query. Thus, a practical approach should not assume that the ground-truth entities in the knowledge graph, which correspond to the mentions in the query, are directly provided by the users.
@@
Conciseness. The retrieved subgraphs should focus on the most relevant and essential nodes and edges, ensuring a clear context for LLMs.
@@
Scalability. An ideal algorithm should scale to large KGs with tens of millions of nodes and edges while maintaining acceptable latency.
These requirements converge on the critical challenge of effectively aligning query text with the underlying structures of a KG. The limitations of existing solutions arise from their intrinsic mechanisms, as summarized in Figure 2. Specifically, (i) KELP Liu et al. (2024) trains a path selection model to identify paths that align with the query text, lacking the plug-and-play usability. (ii) KAPING Baek et al. (2023) and G-retriever He et al. (2024) rely on query text embeddings to retrieve similar triples or connected components in KG, which may introduce noisy information and compromise the conciseness. (iii) KG-GPT Kim et al. (2023a) segments the query text into sub-sentences for precision, but it depends on the LLM to align candidate relations in KG with the sub-sentences, compromising scalability as the number of relations increases. (iv) To compensate for the insufficient alignment capability between query text and KG structures, certain methods inadvertently lead to impractical entity leaks. Both KG-GPT Kim et al. (2023a) and KELP Liu et al. (2024) expand the subgraphs or paths from the exact topic entities. Similarly, G-retriever He et al. (2024) limits the input KG to a 2-hop neighborhood grown from the ground-truth entities of the query.
@@
In this paper, we introduce a novel approach for aligning query text with KG structures. Specifically, we first utilize an LLM to generate a pattern graph that aligns with the query text. Then, to retrieve the best subgraphs from KG that semantically align with the pattern graph, we introduce a novel metric termed Graph Semantic Distance (GSD). Derived from the pairwise matching distance Blumenthal (1953), it quantifies the alignment by summing the semantic distances between corresponding nodes and relations in the pattern graph and the candidate isomorphic subgraphs. For example, in Figure 2, the LLM generates a star-shaped pattern graph aligning with the query. And the highlighted subgraph with the smallest GSD is considered as the best-aligned subgraph in KG.
@@
Figure 3 presents the overview of the proposed Similar Graph Enhanced Retrieval-Augmented Generation (SimGRAG) method.
@@
  • We propose the query-to-pattern and pattern-to-subgraph alignment paradigm, ensuring plug-and-play usability and context conciseness.
  • We define the graph semantic distance and develop an optimized subgraph retrieval algorithm to ensure scalability and avoid entity leaks.
@@ 2 Related Work
The intricate structures of knowledge graphs (KGs) present significant challenges to traditional RAG pipelines, prompting the development of various techniques for graph-based indexing, retrieval, and generation Peng et al. (2024). As depicted in Figure 2, KAPING Baek et al. (2023) retrieves KG triples most relevant to the query directly. KG-GPT Kim et al. (2023a) segments the query and presents LLMs with all candidate relations in the KG for decision-making. KELP Liu et al. (2024) trains a model to encode paths in the KG for selecting relevant paths, although it struggles to scale to structures more complex than 2-hop paths. G-Retriever He et al. (2024) adopts a multi-step approach: it retrieves similar entities and relations, constructs a connected subgraph optimized via the prize-collecting Steiner tree algorithm, and employs a GNN to encode the subgraph for prompt tuning with the LLM.
@@ 3 Preliminaries
A knowledge graph (KG) $\mathcal{G}$ is defined as a set of triples, i.e., $\mathcal{G}=\{(h,r,t)\mid h,t\in\mathcal{V},r\in\mathcal{R}\}$, where $\mathcal{V}$ represents the set of entity nodes and $\mathcal{R}$ denotes the set of relations.
Given a knowledge graph $\mathcal{G}$ and a user query $\mathcal{Q}$, the task of Knowledge Graph Driven Retrieval-Augmented Generation is to generate an answer $\mathcal{A}$ by leveraging both large language models and the retrieved evidence from $\mathcal{G}$. This task is general and encompasses a variety of applications, including but not limited to Knowledge Graph Question Answering (KGQA) and Fact Verification Kim et al. (2023a); Liu et al. (2024).
@@ 4 The SimGRAG Approach
Effectively aligning query text with the intricate structures of KG is a critical challenge for KG-driven RAG approaches. In this section, we introduce a novel strategy that decomposes this alignment task into two distinct phases: query-to-pattern alignment and pattern-to-graph alignment.
@@ 4.1 Query-to-Pattern Alignment
Given a query text $\mathcal{Q}$, we prompt the LLM to generate a pattern graph $\mathcal{P}$ consisting of a set of triples $\{(h_{1},r_{1},t_{1}),(h_{2},r_{2},t_{2}),\dots\}$ that align with the query semantics. To guide the LLM in generating the desired patterns, our prompt first asks for the segmented phrases for each triple before generating all the triples. As shown in Table 11, it also includes several explicit requirements, along with a few input-output examples to facilitate in-context few-shot learning Agarwal et al. (2024).
Such query-to-pattern alignment leverages the inherent understanding and instruction-following capabilities of LLMs. For queries involving up to 3 hops in the FactKG dataset Kim et al. (2023b), Llama 3 70B Dubey et al. (2024) achieves an accuracy of 93% in our experiments. This demonstrates that such alignment can be effectively performed by the LLM without the need for additional training, ensuring plug-and-play usability. Furthermore, as LLMs continue to evolve in both capability and cost-effectiveness, we expect the alignment accuracy to keep improving in the near future.
@@ 4.2 Pattern-to-Subgraph Alignment
The pattern graph $\mathcal{P}$ has a node set $V_{\mathcal{P}}$, while the subgraph $\mathcal{S}$ has a node set $V_{\mathcal{S}}$. We say that $\mathcal{P}$ and $\mathcal{S}$ are isomorphic if there exists a bijective mapping $f:V_{\mathcal{P}}\rightarrow V_{\mathcal{S}}$ s.t. an edge $\langle u,v\rangle$ exists in $\mathcal{P}$ iff the edge $\langle f(u),f(v)\rangle$ exists in $\mathcal{S}$.
@@
After aligning the subgraph structure through graph isomorphism, we proceed to consider the semantic information of the nodes and relations. In traditional text-driven RAG pipelines, computing the distances between the query and document embeddings has proven effective. Similarly, for each entity node $v$ and relation $r$ in both the pattern graph $\mathcal{P}$ and the subgraph $\mathcal{S}$, we obtain the corresponding embedding vectors $z$ as follows:
@@
For a subgraph $\mathcal{S}$ isomorphic to $\mathcal{P}$, the nodes and edges in $\mathcal{S}$ have a one-to-one mapping with those in $\mathcal{P}$. The semantic similarity between the matched nodes or edges can be quantified by computing the L2 distance between their embeddings. Therefore, we use the pairwise matching distance Blumenthal (1953) to derive the following overall graph semantic distance.
@@
Focusing exclusively on isomorphic subgraphs guarantees conciseness. In Section 5, we will provide a detailed discussion on how to efficiently retrieve the top-$k$ isomorphic subgraphs with the smallest GSD in KG.
Furthermore, the joint use of graph isomorphism and semantic similarity effectively reduces noise. In practice, KGs are often noisy, and even semantically similar entities or relations may not always constitute suitable evidence. Figure 4 presents the distance rankings over the 10-million-scale DBpedia for the pattern graph in Figure 2. There are numerous entities related to “Georgian”, but only the entity ranked 112 contributes to the final subgraph. Similarly, for the relation “architecture style”, only the relation ranked 3 is useful. The proposed GSD metric can effectively incorporate somewhat distant entities or relations that still contribute valuable evidence to the overall subgraph, thereby eliminating the need for entity leaks.
@@ 4.3 Generalization to Unknown Entities or Relations
In such cases, we further generalize the Graph Semantic Distance (GSD). Specifically, since unknown entities or relations are ambiguous and difficult to match with corresponding entities or relations in the KG, we exclude them from the GSD computation. Given the isomorphic mapping $f:V_{\mathcal{P}}\rightarrow V_{\mathcal{S}}$ between the pattern graph $\mathcal{P}$ and the KG subgraph $\mathcal{S}$, we generalize GSD to:
@@ 4.4 Verbalized Subgraph-Augmented Generation
Given the top-$k$ subgraphs with the smallest Graph Semantic Distance (GSD) from the KG, we now expect the LLM to generate answers to the original query based on these evidences. To achieve this, we append each retrieved subgraph $\mathcal{S}$ to the query text in the prompt. Each subgraph is verbalized as a set of triples $\{(h_{1},r_{1},t_{1}),(h_{2},r_{2},t_{2}),\dots\}$, as illustrated in Figure 3. Additionally, to facilitate in-context learning, we also include a few example queries with their corresponding subgraphs and expected answers in the prompt. Please refer to Appendix B for more details.
@@ 5.1 Top-$k$ Retrieval Algorithm
5.2 Optimized Retrieval Algorithm
Assume that during the expansion of the $i^{\text{th}}$ edge in $L$, the partial mapping from the pattern graph $\mathcal{P}$ to the knowledge graph $\mathcal{G}$ is represented by $f$. Suppose there exists an isomorphic mapping $f^{\prime}$ that can be completed by future expansion, resulting in a subgraph $\mathcal{S}$ with $GSD(\mathcal{P},\mathcal{S})$. It can be decomposed into the following four terms, where $L[1:i]$ denotes the first $i-1$ triples in $L$ and $L[i:]$ denotes the remaining triples.
For Equations (6) and (8), notice that
Combining Equations (4), (9), and (10), we have
By combining the pruned and greedy expansion strategies, the optimized algorithm is guaranteed to produce the same results as the top-$k$ retrieval algorithm without any loss in solution quality. Experiments in Section 6.6 demonstrate that the optimized algorithm significantly accelerates retrieval.

### v2

@@ Abstract
Recent advancements in large language models (LLMs) have shown impressive versatility across various tasks. To eliminate their hallucinations, retrieval-augmented generation (RAG) has emerged as a powerful approach, leveraging external knowledge sources like knowledge graphs (KGs). In this paper, we study the task of KG-driven RAG and propose a novel Similar Graph Enhanced Retrieval-Augmented Generation (SimGRAG) method. It effectively addresses the challenge of aligning query texts and KG structures through a two-stage process: (1) query-to-pattern, which uses an LLM to transform queries into a desired graph pattern, and (2) pattern-to-subgraph, which quantifies the alignment between the pattern and candidate subgraphs using a graph semantic distance (GSD) metric. We also develop an optimized retrieval algorithm that efficiently identifies the top-$k$ subgraphs within 1-second on a 10-million-scale KG. Extensive experiments show that SimGRAG outperforms state-of-the-art KG-driven RAG methods in both question answering and fact verification. Our code is available at https://github.com/YZ-Cai/SimGRAG.
@@ 1 Introduction
Pre-trained large language models (LLMs) are popular for diverse applications due to their generality and flexibility Zhao et al. (2023); Minaee et al. (2024); Wang et al. (2024a). To avoid the hallucinations or outdated knowledge of LLMs Zhang et al. (2023); Baek et al. (2023), Retrieval-Augmented Generation (RAG) Zhao et al. (2024); Gao et al. (2023) integrates LLMs with external knowledge sources to produce grounded outputs, where knowledge graphs (KGs) Ji et al. (2022) have emerged as a valuable option Peng et al. (2024).
For many KG-driven tasks, their KG schemas align with human cognition and can be read by humans. In other words, a non-specialist can describe the knowledge using an intuitive graph structure. In this paper, we follow existing KG-driven RAG methods Baek et al. (2023); Kim et al. (2023a); Liu et al. (2024) and focus on such human-understandable KGs to enable the mimicking of human reasoning. As shown in Figure 1, an ideal approach should address the following features.
@@
Plug-and-Play on Human-Understandable KGs. To fully leverage the inherent generalization power of LLMs, an ideal approach should be easily deployable without additional training or fine-tuning for KGs that align with human cognition and can be interpreted by LLMs. Otherwise, training a smaller and task-specific model on such KGs would be a more cost-effective alternative.
Avoidance of Requiring Oracle Entities. In real applications, users might not always know the precise entity IDs related to their queries. Thus, it would be better if a method naturally does not require users to specify the oracle entities.
@@
Context Conciseness. The retrieved subgraphs should focus on the most relevant and essential nodes and edges, ensuring clear contexts for LLMs.
@@
Retrieval Scalability. An ideal algorithm should scale to large KGs with tens of millions of nodes and edges while maintaining acceptable latency.
Existing approaches typically follow a paradigm of retrieving subgraphs from the KG and feeding them into LLMs to generate the final response. The critical challenge lies in effectively aligning query texts with the structural knowledge encoded in KGs. Figure 2 summarizes different mechanisms of existing approaches. Specifically, (i) KAPING Baek et al. (2023) employs query text to directly retrieve isolated triples using their semantic embedding similarity, which struggles with multi-hop queries as the query embedding captures excessive information. (ii) G-retriever He et al. (2024) uses query text embeddings to retrieve similar entities and relations, then extracts a connected components in KG, which potentially cannot guarantee the best conciseness of the retrieved subgraphs. (iii) KG-GPT Kim et al. (2023a) segments the query into sub-sentences but depends on the LLM to decide relations in KG that can match each sub-sentence, compromising scalability as the number of candidate relations increases. (iv) KELP Liu et al. (2024) trains a path selection model to identify paths that align with the query text, lacking the plug-and-play usability even on human-understandable KGs.
@@
In this paper, we introduce a novel approach, Similar Graph Enhanced Retrieval-Augmented Generation (SimGRAG) method, for aligning query text with KG structures. Figure 3 presents the overview with 3 steps. (1) Query-to-Pattern Alignment. We utilize an LLM to generate a pattern graph that aligns with the query text. (2) Pattern-to-Subgraph Alignment. To retrieve the best subgraphs from KG that semantically align with the generated pattern graph, we introduce a novel metric termed Graph Semantic Distance (GSD). It quantifies the alignment by summing the semantic distances between corresponding nodes and relations in the pattern graph and the candidate isomorphic subgraphs. For example, in Figure 2, the LLM generates a star-shaped pattern graph aligning with the query. And the highlighted subgraph with the smallest GSD is considered as the best-aligned subgraph in KG. (3) Verbalized Subgraph Augmented Generation. Finally, the query and the retrieved subgraphs are passed to an LLM to generate the answer.
@@
  • We propose the query-to-pattern and pattern-to-subgraph alignment paradigm, ensuring the plug-and-play usability on human-understandable KGs and the context conciseness for LLMs.
  • We define the graph semantic distance and develop an optimized subgraph retrieval algorithm to avoid requiring oracle entities and ensure retrieval scalability on million-scale KGs.
@@ 2 Related Work
Graph Retrieval-Augmented Generation.
Graph Retrieval-Augmented Generation (GraphRAG) integrates graphs into RAG pipelines, which can be categorized into 10 domains, including knowledge graph (KG), document graph and so on Han et al. (2024). GraphRAG methods may use existing graphs or construct graphs from other data source, such as building a knowledge graph (KG) from documents Choubey et al. (2024). We focus on the KG-driven RAG scenario, which utilizes existing manually constructed KGs that used for retrieval in the RAG pipeline, as detailed as follows.
@@
The intricate structures of knowledge graphs (KGs) present significant challenges to traditional RAG pipelines, prompting the development of various techniques for graph-based indexing, retrieval, and generation Peng et al. (2024). As depicted in Figure 2, KAPING Baek et al. (2023) retrieves KG triples most relevant to the query directly. KG-GPT Kim et al. (2023a) segments the query and presents LLMs with all candidate relations in the KG for decision-making. KELP Liu et al. (2024) trains a model to encode paths in the KG for selecting relevant paths, although it struggles to scale to structures more complex than 2-hop paths. G-Retriever He et al. (2024) first retrieves similar entities and relations, then constructs a connected subgraph optimized via the prize-collecting Steiner tree algorithm, and employs a GNN to encode the subgraph for prompt tuning with the LLM.
@@ 3 Preliminaries
A knowledge graph (KG) $\mathcal{G}$ is defined as a set of triples, i.e., $\mathcal{G}=\{(h,r,t)\mid h,t\in\mathcal{V},r\in\mathcal{R}\}$, where $\mathcal{V}$ represents the set of entity nodes and $\mathcal{R}$ denotes the set of relations. Given a knowledge graph $\mathcal{G}$ and a user query $\mathcal{Q}$, the task of Knowledge Graph Driven Retrieval-Augmented Generation is to generate an answer $\mathcal{A}$ by leveraging both large language models and the retrieved evidence from $\mathcal{G}$. This task is general and encompasses a variety of applications, including but not limited to Knowledge Graph Question Answering (KGQA) and Fact Verification Kim et al. (2023a); Liu et al. (2024).
@@ 4 The SimGRAG Approach
Effectively aligning query text with the KG structures is a critical challenge. In this section, we introduce a novel strategy that decomposes this alignment task into two distinct phases: query-to-pattern alignment and pattern-to-graph alignment.
@@ 4.1 Query-to-Pattern Alignment
Given a query text $\mathcal{Q}$, we prompt the LLM to generate a pattern graph $\mathcal{P}$ consisting of a set of triples $\{(h_{1},r_{1},t_{1}),(h_{2},r_{2},t_{2}),\dots\}$ that align with the query semantics. We expect the LLM to interpret the user query thoughtfully, but we do not expect it to produce the exact same entities or relations appeared in the KG.
To guide the LLM in generating the desired patterns, our prompt first asks for the segmented phrases for each triple before generating all the triples. As shown in Table 17, it also includes a few explicit requirements. To facilitate in-context few-shot learning Agarwal et al. (2024), we further manually construct a few examples (typically 12-shots) based on the characteristics of each KG, guiding the LLM to generate desired patterns.
Such query-to-pattern alignment leverages the inherent understanding and instruction-following capabilities of LLMs. Based on our experiments detailed in Section 6, the accuracy of the alignment can be defined as the proportion of queries that conform to the expected pattern under manual verification. For queries involving up to 3 hops in the MetaQA Zhang et al. (2018) and FactKG Kim et al. (2023b) datasets, Llama 3 70B Dubey et al. (2024) achieves the accuracies of 98% and 93%, respectively. Thus, on KGs following human cognition which can be understood by humans, such alignment could be effectively performed by the LLM without the need for additional training, ensuring plug-and-play usability. But for certain KGs with specialized structures, it may be inevitable to further fine-tune the LLMs for mimicing domain-specific specialists, as discussed in Section Limitations.
@@ 4.2 Pattern-to-Subgraph Alignment
The pattern graph $\mathcal{P}$ has a node set $V_{\mathcal{P}}$, while the subgraph $\mathcal{S}$ has a node set $V_{\mathcal{S}}$. We say that $\mathcal{P}$ and $\mathcal{S}$ are isomorphic if there exists a bijective mapping $f:V_{\mathcal{P}}\rightarrow V_{\mathcal{S}}$ s.t. an edge $\langle u,v\rangle$ exists in $\mathcal{P}$ if and only if the edge $\langle f(u),f(v)\rangle$ exists in $\mathcal{S}$.
@@
After aligning the subgraph structure through graph isomorphism, we proceed to consider the semantic information of the nodes and relations. Similar to traditional text-driven RAG pipelines, for each entity node $v$ and relation $r$ in both the pattern graph $\mathcal{P}$ and the subgraph $\mathcal{S}$, we obtain the corresponding embedding vectors $z$ as follows:
@@
In this paper, we use the Nomic embedding model Nussbaum et al. (2024), which generates 768-dim semantic embeddings for nodes and relations.
For a subgraph $\mathcal{S}$ isomorphic to $\mathcal{P}$, the nodes and edges in $\mathcal{S}$ have a one-to-one mapping with those in $\mathcal{P}$. By computing the L2 distance between their embeddings, we use the pairwise matching distance Blumenthal (1953) to derive the following overall graph semantic distance.
@@
Focusing exclusively on isomorphic subgraphs guarantees conciseness. Section 5 will provides algorithms to efficiently retrieve the top-$k$ isomorphic subgraphs with the smallest GSD in KG.
Furthermore, the joint use of graph isomorphism and semantic similarity effectively reduces noise. In practice, KGs are often noisy, and even semantically similar entities or relations may not always constitute suitable evidence. Figure 4 presents the distance rankings over the 10-million-scale DBpedia for the pattern graph in Figure 2. There are numerous entities related to “Georgian”, but only the entity ranked 112 contributes to the final subgraph. Similarly, for the relation “architecture style”, only the relation ranked 3 is useful. The proposed GSD metric can effectively incorporate somewhat distant entities or relations that still contribute valuable evidence to the overall subgraph, thereby eliminating the need for oracle entities.
@@ 4.3 Generalization to Unknown Entities or Relations
In such cases, we further generalize the Graph Semantic Distance (GSD). Specifically, since the unknown entities or relations are ambiguous and difficult to match with corresponding entities or relations in the KG, we exclude them from the GSD computation. Given the isomorphic mapping $f:V_{\mathcal{P}}\rightarrow V_{\mathcal{S}}$ between the pattern graph $\mathcal{P}$ and the KG subgraph $\mathcal{S}$, we generalize GSD to:
@@ 4.4 Verbalized Subgraph-Augmented Generation
Given the top-$k$ subgraphs with the smallest Graph Semantic Distance (GSD) from the KG, we now expect the LLM to generate answers to the original query based on these evidences. To achieve this, we append each retrieved subgraph $\mathcal{S}$ to the query text in the prompt. Each subgraph is verbalized as a set of triples $\{(h_{1},r_{1},t_{1}),(h_{2},r_{2},t_{2}),\dots\}$, as illustrated in Figure 3. Additionally, to facilitate in-context learning, we also manually curate a few example queries (typically 12-shots) with their corresponding subgraphs and expected answers in the prompt. Please refer to Appendix B for details.
@@ 21                         Expand($i+1,f\cup\{t_{\mathcal{P}}:t_{\mathcal{G}}\}$);
5.2 Optimized Retrieval Algorithm
Assume that during the expansion of the $i^{\text{th}}$ edge in $L$, the partial mapping from $\mathcal{P}$ to the knowledge graph $\mathcal{G}$ is represented by $f$. Suppose there exists an isomorphic mapping $f^{\prime}$ that can be completed by future expansion, resulting in a subgraph $\mathcal{S}$ with $GSD(\mathcal{P},\mathcal{S})$. It can be decomposed into four terms, where $L[1:i]$ denotes the first $i-1$ triples in $L$ and $L[i:]$ denotes the remaining triples.
For Equations (6) and (8), notice that
Combining Equations (4), (9), and (10), we have
By combining the pruned and greedy expansion strategies, the optimized algorithm is guaranteed to produce the same results as the top-$k$ retrieval algorithm without any loss in solution quality. The experiments in Section 6.6 show that the optimized algorithm significantly accelerates retrieval.